OpenStudy (anonymous):

solve for t. 2500=500e^0D6t

OpenStudy (anonymous):

I would advise Log_5 \[Log_5 \]Both sides

OpenStudy (anonymous):

hey i got this give me a sec to type

OpenStudy (anonymous):

\[D= \sqrt[6]{5 t5}/t, - \sqrt[6]{5t5}/t\]

OpenStudy (anonymous):

A pic i made real fast

OpenStudy (anonymous):

plz say good answer if i helped

OpenStudy (anonymous):

you did thank you!

OpenStudy (anonymous):

Plz click good answer

OpenStudy (anonymous):

How i do supper man?

OpenStudy (anonymous):

2500=500e^6t 2500/500 = e^ 6t 5 =e^ 6t take ln both sides ln 5 = ln (e^ 6t) ln 5 = 6t ln e ln 5 =6t ln 5/ 6 =t 1.61/6 =t 0.27 =t or, t = 0.27